OpenStudy (kamille):

I know this is quite easy, but I need to be reminded how to solve it. The pool was filled by 5 taps in 42 minutes. How much time it would take to fill the pool if there is 6 taps? By the way, the answer is 35minutes.

OpenStudy (anonymous):

You need to solve 5*42 = 6*T where T is the time taken to fill the pool if there are 6 taps. Rearranging this you get T = 210/6 = 35.

OpenStudy (kamille):

oh, it was really easy. I have always thought I need to make a proportion like 5 taps - 42 minutes 6 taps - x minutes but I have always get wrong answer! Thank you!!!